OpenStudy (anonymous):

How do you solve 0 = -3/5x - 3 ?

OpenStudy (anonymous):

\[0 = \frac{3}{5}x - 3\]\[3 = \frac{3}{5}x\]\[15 = 3x\]\[5 = x\]

OpenStudy (anonymous):

Would it be -5, since the fraction is negative?

OpenStudy (anonymous):

Oops. Yes. :D

OpenStudy (anonymous):

Thanks :)